AB INITIO STUDY OF THE TRICHLORINE RADICAL, Cl_{3}

\underline{A. L. KALEDIN}, W. G. LAWRENCE, M. C. HEAVEN, Q. CUI, J. E. STEVENS, and K. MO-ROKUMA, Cherry L. Emerson Center for Scientific Computation and Department of Chemistry, Emory University, Atlanta, GA 30322, U.S.A.

We report a rigorous ab initio study of the ground and low-lying excited-state potential energy surfaces (PES) of the Cl_{3} radical at CASSCF, CASPT2, and internally contracted (IC) MRSDCI levels of theory with the Dunning’s avdz and avtz basis sets. ( There is a previous study ^{a} which examined only restricted areas of the PES. ) The ground state PES has two true minima, both of which are van der Waals complexes between Cl and Cl_{2}. The linear asymmetric minimum is of ^{2}Π symmetry, and the bent asymmetric minimum is ^{2}A′. At the ICMRSDCI/avdz level of theory with counterpoise correction for energy, the Jacobi coordinates for the linear minimum are: r=3.90 au, R=8.47 au, D_{e}(Cl_{2}(X)-Cl)= 250 cm^{−1}, and those of the bent minimum are: r=3.90 au, R=6.85 au, γ=66.47^{o}, D_{e}(Cl_{2}(X)-Cl)=230 cm^{−1}. Addition of spin-orbit interaction as a perturbation predicts that the global minimum is linear ^{2}Π_{3/2} stabilized by 280 cm^{−1}whereas bent ^{2}E_{1/2} is stabilized by only 136 cm^{−1}. Excited valence states exhibit only one strongly bound minimum: ^{2}Π_{g} with the bond length of 4.67 au at the ICMRS-DCI/avdz level of theory. This state is located 1.1 eV above the van der Waals minima and exhibits a strong (1 au) dipole transition moment to the below-lying ^{2}Π_{u} state. It is bound by approximately 5,000 cm^{−1}with respect to Cl_{2}(^{3}Π_{u})-Cl asymptote. This minimum is due to the avoided crossing between two valence bond structures: Cl-Cl Cl and Cl Cl-Cl. Attempts are made in connection to previous experimental and ab initio studies to explain the on-going hypothesis of a long-lived ground state intermediate formed by the reaction: Cl_{2} + Cl ⇔ Cl_{3}. According to our ground state calculations this postulate is strongly undermined. In addition, we attempt to explain a 470 nm long-lived laser-induced fluorescence observed after isolating Cl and Cl_{2} in Ar matrix. Our calculations do not show any evidence that the trichlorine radical is responsible for this fluorescence. However, the possibilities of Cl_{3} playing some role in this process are examined.

aA. B. Sannigrahi and Sigrid D. Peyerimhoff, Int. J. Quantum Chem. 30, 413-420 (1986).